the question is:
solve this system.
x-y=5
x2+y2=25

Answer by oscargut(2103)

x-y=5
x2+y2=25
from the first eq we have: y=x-5
using 2nd eq x^2+(x-5)^2=25
2x^2-10x+25=25
2x^2-10x=0
2x(x-5)=0
x=0 or x=5
y=-5 or y=0
Solutions are: (0,-5) and (5,0)